A friend taking this class asks you to explain why, when solving for values or doing problems with given numbers, ΔG° is rarely zero, but ΔG is often zero. Explain using energy diagrams (feel free to model a hypothetical reaction A  B, or anything else you like).
Answer #1

We know that,

Delta G° = - n*F*E0

Where n is the number of electrons transferred, F is Faraday constant and E0 is the standard cell potential.

As Standard potential of a cell cannot be zero so, Delta G° cannot be zero

Again,

Delta G = -n* F* E

E = E0 - RT * lnA/nF

Where A is reaction quotient. If it can be adjusted in such a way that

E0 = RT*lnA / nF

In this case, E = 0 and hence Delta G will be 0.